# Nondeterminism-Aware Optimistic Verification for Floating-Point Neural Networks

### v2

@@ 6.4. Single-operator verification
Routing policy. At the single operator $v^{*}$ (inputs and attributes already committed), the challenger chooses one of two verification paths. Routing policy: off-chain the challenger recomputes a reference $y_{v^{*}}^{\mathrm{ref}}$ and compares the elementwise error $|y_{v^{*}}^{P}-y_{v^{*}}^{\mathrm{ref}}|$ to the theoretical cap $\tau^{\text{theo}}_{v^{*}}$. If any element exceeds $\tau^{\text{theo}}_{v^{*}}$, it invokes path (i); otherwise, it invokes path (ii) to apply the tighter empirical thresholds.
(i) Theoretical-bound verification. The challenger requests the proposer to submit a verifiable proof (through opVM/zkVM) that, given the committed inputs and operator metadata: (i) a canonical reference execution computes $y_{v^{*}}^{\mathrm{ref}}$ under IEEE-754 FP32 semantics; (ii) the operator-wise theoretical bound $\tau^{\text{theo}}_{v^{*}}$ is computed; (iii) element-wise, $|y_{v^{*}}^{P}-y_{v^{*}}^{\mathrm{ref}}|\leq\tau^{\text{theo}}_{v^{*}}$. The on-chain verifier checks the single proof. If it passes the verification, the computation is accepted; otherwise, the proposer loses the dispute.
@@
6.5. Economic soundness and incentives.
While the dispute game ensures correctness at the technical level, it can be combined with established economic mechanisms that allocate rewards and penalties to achieve incentive compatibility for proposers, challengers, and committee members, while also discouraging collusion. Such mechanisms have been extensively developed in other accountability frameworks for off-chain execution, including incentive-secure proof-of-learning protocols (Zhao et al., 2025), sampling-and-penalty schemes (Zhang et al., 2024), and watchtower-style networks (Sheng et al., 2024). We directly use these mature mechanisms to support our protocol: honest execution remains individually rational, voluntary challenges are properly incentivized, and committee participation under randomized audits is sustainable.
@@ 7. Implementation and evaluation
We implement NAO as (i) a PyTorch runtime that instruments graphs, computes bounds, splits/executes subgraphs, proceeds dispute games, and emits/verifies commitments, and (ii) a set of contracts on Ethereum Holesky orchestrating commitments, dispute state, per-round bonds. The runtime adds no custom kernels and runs unmodified models.
@@ 7.1. Experimental setup
Hardware and software configuration. Empirical error percentile threshold calibration uses four GPUs (NVIDIA RTX 4090, A100, H100, RTX 6000). We use PyTorch with FP32 forwards and FP64 for bound arithmetic. We set: fixed RNG seeds, fast-math disabled, CUDA/cuDNN library determinism, TF32/benchmark disabled, deterministic cuBLAS via CUBLAS_WORKSPACE_CONFIG. Contracts run on Ethereum Holesky testnet; we report on-chain costs in k gas. We ignore network latency during interactions.
Models and datasets. We evaluate ResNet-152 (He et al., 2016) on ImageNet (Deng et al., 2009), BERT-large (Devlin et al., 2019) on DBpedia (Auer et al., 2007), Qwen3-8B (Team, 2025) on C4 (Raffel et al., 2019) (next-token prediction), and Stable Diffusion v1-5 (Rombach et al., 2022). PyTorch graph sizes range from $1k$ to $5k$ operators.
@@ 7.2. System microbenchmarks and error analysis
Figure 7. Microbenchmarks on BERT-large dispute game: vary split size NN and dispute round index. Left: avg dispute rounds decrease roughly as $O(\log_{N}|V|)$. Mid-left: avg dispute time drops sharply from small $N$ and plateaus for $N\!\geq\!8$. Mid-right: avg Merkle proof checks for weights/graph fall as NN grows. Right: per-round substep time (proposer partition vs. challenger re-execute plus selection) measured across eight perturbed operators through the model.:
@@
Implications for verification. First, search-time checks are one–three orders of magnitude tighter than worst-case theory for transformers; evasion would require perturbations that are consistently sub-micro across many operators, aligning with the non-accumulation results in Sec. 5. Second, leaf-only verification under theoretical bounds leaves more slack: especially for LLM stacks, motivating the committee option at the leaf when tighter adjudication is desired. We subsequently confirm these implications in the attack experiments, where empirical error percentile thresholds yield $0\%$ ASR across models while worst-case bounds allow a small attack window (Sec. 5.5).
Dispute game statistics and scaling with $N$. Fig. 7 varies the partition size $N$. Avg. Dispute Rounds: increasing $N$ reduces rounds from $\sim\!11$ at N=2N{=}2 to $\sim\!3$ at $N{\geq}12$. Avg. Dispute Time: dispute time drops sharply from N=2N{=}2 to $N{\approx}6$ and then flattens; the first round dominates latency because it handles the largest subgraph. On-chain work: Merkle proof checks (graph and weights) shrink monotonically with $N$ (roughly $2{\times}$$4{\times}$ reduction between N=2N{=}2 and $N{\geq}8$). Per-round substep latency: the rightmost panel shows the time for the proposer’s partition and the challenger’s selection at dispute round index $i$ within an $N=4$-way partition; both decay with $i$. Boxplot information is collected through eight dispute games across eight different perturbed operators through the model. Guideline and insight: $N\!\in\![8,12]$ minimizes wall-clock and on-chain work without creating oversized proofs.
@@ 7.3. End-to-end cost and on-chain footprint
Table 3. Forward and dispute costs across models ($N=2$). DCR is the challenger FLOP range to reach and adjudicate the leaf; Cost Ratio is DCR/forward FLOPs. Gas sums all dispute message interactions.
@@
Deterministic execution overhead. The deterministic settings introduced in Sec. 7.1 is also set during the optimistic execution phase (Phase 1 in Sec. 3.2) to ensure the system functions correctly. We test the overhead of these settings by running Qwen-8B model on 100 WikiText-103 (Merity et al., 2017) input data. The experiments show that the software deterministic settings introduce a negligible overhead of  0.3% in terms of latency compared to the non-deterministic settings.
@@
Where NAO fits. When users accept tolerance-aware correctness, NAO offers orders-of-magnitude lower end-to-end latency and memory usage than zk proving for LLM-scale inference while preserving native kernels and heterogeneity. zk remains compelling when public verifiability with model secrecy is paramount; NAO is compelling when floating-point semantics and production performance are constraints.
@@ 8.1. Verifiable Machine Learning
Work on verifiable ML splits into cryptographic validity proofs and optimistic approaches. Zero-knowledge systems compile models to finite-field circuits and attach succinct proofs (e.g., SafetyNets (Ghodsi et al., 2017), zkML (Chen et al., 2024), zkGPT (Qu et al., 2025), zkllm (Sun et al., 2024)), but still incurring substantial proving time/memory and often requiring quantization or restricted operators; TEE-assisted designs reduce proof cost but reintroduce trust in hardware vendors and exists performance degradation due to limited storage and compute capacity (Tramer and Boneh, [n. d.]; Moon et al., 2025; Bai et al., 2025; Niu et al., 2022). Optimistic protocols (Agatha (Zheng et al., 2021), opML (Conway et al., 2024)) and refereed delegation (Verde (Arun et al., 2025)) localize disputes on computation graphs and arbitrate on chain, with practicality often hinging on replication or enforcing bitwise-reproducible operators (RepOps) across heterogeneous accelerators; recent systems generalize dispute games to Turing-complete programs and explore incentive mechanisms such as Proof-of-Sampling (Mirkin et al., 2025; Zhang et al., 2024). NAO takes a complementary tack: rather than enforcing bitwise equality, it verifies results up to principled, operator-level tolerances that acknowledge nondeterminism. Unlike approaches that determinize execution (e.g., high-precision rounding logs for verifiable training or RepOps libraries (Srivastava et al., 2024; Arun et al., 2025)), NAO aims to preserve native GPU kernels and heterogeneity while still providing portable accountability.
@@ 8.2. Floating-Point Rounding Errors in Neural Networks
A growing body of work shows that ignoring floating-point effects renders verification unsound in practice: real-number proofs can certify networks that fail under IEEE-754 execution, and deployed stacks exhibit nondeterminism from non-associative reductions, kernel fusion, and scheduler variation (Jia and Rinard, 2021; Szász et al., 2025). Measurement tools such as FPRev reverse-engineer accumulation orders across software/hardware stacks(Xie et al., 2025). In response, some systems try to eliminate nondeterminism, e.g., reproducible operator libraries or training with high-precision states and logged rounding decisions to enable exact replay across GPUs, but at non-trivial performance and engineering cost (Arun et al., 2025; Srivastava et al., 2024).
9. Discussion and Conclusion
We introduce NAO, a tolerance-aware optimistic protocol that makes neural network computation verifiable on heterogeneous accelerators without relying on bitwise determinism or trusted hardware. NAO verifies results up to operator-specific acceptance regions by combining portable IEEE-754 bounds with tight empirical error percentile thresholds, and it localizes disputes to a single operator via a Merkle-anchored dispute game followed by a low-cost leaf check. This design preserves native kernels and performance while providing vendor-agnostic accountability; our implementation and attack study demonstrate the robustness of the tolerance-aware approach.
NAO naturally handles multi-step workloads (autoregressive decoding, diffusion sampling, and training) by layering time over the dispute game. Concretely, we commit to a temporal Merkle tree of step states (e.g., token logits, parameter checkpoints), bisect first across time to the earliest offending step, then within that step across the operator DAG to the leaf adjudication. This 2D partition yields prefix finality: early tokens/steps can finalize even while later steps remain subject to challenge. For training, we anchor checkpoints and optimizer state, verify per-update deltas, and permit windowed challenges over mini-batches. A minor limitation is calibration churn: onboarding new devices or kernels requires efficiently updating empirical thresholds and validating them. Extending NAO so that multi-step settings work by generalizing temporal bisection/$N$ way partition with windowed mini-batch challenges and streamlining threshold calibration is ongoing work.

### v3

@@ 5.4. Single-operator verification
Routing policy. At the single operator $v^{*}$ (inputs and attributes already committed), the challenger chooses one of two verification paths. Routing policy: the challenger recomputes a reference $y_{v^{*}}^{\mathrm{ref}}$ and compares the elementwise error $|y_{v^{*}}^{P}-y_{v^{*}}^{\mathrm{ref}}|$ to the theoretical cap $\tau^{\text{theo}}_{v^{*}}$. If any element exceeds $\tau^{\text{theo}}_{v^{*}}$, it invokes path (i); otherwise, it invokes path (ii) to apply the tighter empirical thresholds.
(i) Theoretical-bound verification. The challenger requests the proposer to submit a verifiable proof that, given the committed inputs and operator metadata: (i) a canonical reference execution computes $y_{v^{*}}^{\mathrm{ref}}$ under IEEE-754 FP32 semantics; (ii) the operator-wise theoretical bound $\tau^{\text{theo}}_{v^{*}}$ is computed; (iii) element-wise, $|y_{v^{*}}^{P}-y_{v^{*}}^{\mathrm{ref}}|\leq\tau^{\text{theo}}_{v^{*}}$. The coordinator verifies the single proof. If it passes the verification, the computation is accepted; otherwise, the proposer loses the dispute.
@@
5.5. Economic soundness and incentives
Complementing the technical correctness of the dispute game, we design an economic mechanism on a blockchain coordinator to ensure incentive compatibility among rational proposers, challengers, and committee members. Adapting the fee-and-deposit pattern from optimistic accountability systems (kalodner2018arbitrum; zhang2024proof; sheng2024proof; zhao2025PoL), TAO requires proposers and challengers to stake deposits $D_{p}$ and $D_{ch}$. A smart contract coordinator escrows these funds and automatically slashes the losing party with $S_{\mathrm{slash}}$ upon adjudication. Meanwhile, committee members receive fixed compensation for participating in the arbitration process, regardless of the challenge outcome.
Two detection channels: voluntary challenges vs. randomized audits. TAO supports two mutually exclusive supervisory channels per claim: (i) a permissionless voluntary challenge with probability $\phi_{ch}$, and (ii) a randomized audit triggered with probability $\phi$, where $\phi+\phi_{ch}\leq 1$. The costs for randomized audits are collected from users as service fees to ensure baseline security. In contrast, the cost of a voluntary challenge is borne by the challenger, where the reward for a successful challenge is set to strictly exceed this cost. Since voluntary challenges provide security redundancy without increasing the systematic fee burden, there is no rationale to exclude this cost-free oversight layer.
Let $\epsilon_{1}$ be the false negative rate (fraud missed due to tolerance allowances) and $\epsilon_{2}$ the false positive rate (wrong detection), where $\epsilon_{1},\epsilon_{2}\in[0,1)$. Then the probability that a cheating proposer is detected and penalized is
(16) $d(\phi,\phi_{ch},\epsilon_{1})\;=\;(\phi+\phi_{ch})\,(1-\epsilon_{1}).$
Fig. 6 summarizes interactions in this mechanism.
Figure 6. Mechanism overview (rewards, penalties, and two detection channels).
Incentive constraints. We choose penalties and deposits so that (a) honest execution dominates cheap cheating, (b) challenging is profitable when fraud exists but spamming is not, and (c) committee participation is sustainable.
Proposer. The proposer faces three possible strategies: honest computation ($h$), cheap cheating ($c_{1}$, e.g., using a smaller model), and targeted cheating ($c_{2}$, e.g., adversarial perturbation injection), with costs $C_{p}$, $C^{\prime}_{p}$, and $C^{\prime\prime}_{p}$. Let $R_{p}$ be the task reward. Expected payoffs are
(17) $u_{p}(h)$ $=R_{p}-C_{p}-\epsilon_{2}S_{\mathrm{slash}},$
(18) $u_{p}(c_{1})$ $=R_{p}-C^{\prime}_{p}-d(\phi,\phi_{ch},\epsilon_{1})\,S_{\mathrm{slash}},$
(19) $u_{p}(c_{2})$ $=R_{p}-C^{\prime\prime}_{p}.$
Honest execution is individually rational when $R_{p}\geq C_{p}+\epsilon_{2}S_{\mathrm{slash}}$. Moreover, to deter cheap cheating it suffices to make honesty strictly better than $c_{1}$:
(20) $u_{p}(h)>u_{p}(c_{1})\quad\Longleftrightarrow\quad S_{\mathrm{slash}}\;>\;\frac{C_{p}-C^{\prime}_{p}}{d(\phi,\phi_{ch},\epsilon_{1})-\epsilon_{2}},$
assuming $d(\phi,\phi_{ch},\epsilon_{1})>\epsilon_{2}$. Note that for $c_{2}$, empirical evidence shows that effective targeted cheating requires $C^{\prime\prime}_{p}\gg R_{p}$, which implies $u_{p}(c_{2})\leq 0$.
Voluntary challenger. Challengers freely select their verification strategy, which practically involves task re-execution and leaf verification at a total cost denoted by $C_{ch}$. A winning challenger receives a share $\alpha_{ch}\,S_{\mathrm{slash}}$ (where $\alpha_{ch}\in(0,1]$). The expected payoffs given the proposer is guilty/clean are
(21) $u_{ch}\big|\text{guilty}$ $=(1-\epsilon_{1})\,\alpha_{ch}\,S_{\mathrm{slash}}-C_{ch},$
(22) $u_{ch}\big|\text{clean}$ $=-C_{ch}-(1-\epsilon_{2})\,D_{ch}.$
We have $u_{ch}|\text{clean}\leq 0$ to deter spam. Conversely, a sufficient condition for honest challenges to be profitable is
(23) $S_{\mathrm{slash}}>\frac{C_{ch}}{\alpha_{ch}\,(1-\epsilon_{1})}.$
Audit committee. In a randomized audit, the coordinator samples $n$ committee members to adjudicate the claim. A sampled member $i$ incurs cost $C_{a}$. The coordinator pays $F_{i}$ (user service fee) to the committee member if the claim is ruled clean, and pays a share $\alpha_{cm}\,S_{\mathrm{slash}}/n$ (where $\alpha_{cm}\in(0,1]$ and $\alpha_{cm}+\alpha_{ch}\leq 1$) if guilt is found. Ex-post payoffs are
(24) $u_{i}^{\text{cm}}\big|\text{ruled guilty}$ $=\tfrac{\alpha_{cm}\,S_{\mathrm{slash}}}{n}-C_{a},$
(25) $u_{i}^{\text{cm}}\big|\text{ruled clean}$ $=F_{i}-C_{a},$
which are strictly positive whenever $\frac{\alpha_{cm}\,S_{\mathrm{slash}}}{n}>C_{a},\;F_{i}>C_{a}$.
Nonempty feasible region for $S_{\mathrm{slash}}$. Let $\mathcal{L}\triangleq\max\{\mathcal{L}_{1},\allowbreak\mathcal{L}_{2},\allowbreak\mathcal{L}_{3}\}$. where $\mathcal{L}_{1}=\frac{C_{p}-C^{\prime}_{p}}{d(\phi,\phi_{ch},\epsilon_{1})-\epsilon_{2}}$, $\mathcal{L}_{2}=\frac{C_{ch}}{\alpha_{ch}\,(1-\epsilon_{1})}$, and $\mathcal{L}_{3}=\frac{n\,C_{a}}{\alpha_{cm}}$. Assuming finite costs, and choosing knobs so that $d(\phi,\phi_{ch},\epsilon_{1})>\epsilon_{2}$, any $S_{\mathrm{slash}}\in(\mathcal{L},D_{p}]$ satisfies all constraints and non-empty by setting $D_{p}>\mathcal{L}$.
@@ 6. Implementation and evaluation
We implement TAO as (i) a PyTorch runtime that instruments graphs, computes bounds, splits/executes subgraphs, proceeds dispute games, and emits/verifies commitments, and (ii) a set of smart contracts on Ethereum Holesky as a coordinator backend orchestrating commitments, dispute state, per-round bonds, and cost measurements. The runtime adds no custom kernels and runs unmodified models.
@@ 6.1. Experimental setup
Hardware and software configuration. Empirical error percentile threshold calibration uses four GPUs (NVIDIA RTX 4090, A100, H100, RTX 6000). We use PyTorch with FP32 forwards and FP64 for bound arithmetic. We set: fixed RNG seeds, fast-math disabled, CUDA/cuDNN library determinism, TF32/benchmark disabled, deterministic cuBLAS via CUBLAS_WORKSPACE_CONFIG. Contracts run on Ethereum Holesky testnet; we report on-chain coordinator computation costs in k gas. We ignore network latency during interactions.
Models and datasets. We evaluate ResNet-152 (he2016deep) on ImageNet (deng2009imagenet), BERT-large (devlin2019bert) on DBpedia (auer2007dbpedia), Qwen3-8B (qwen3technicalreport) on C4 (2019t5) (next-token prediction), and Stable Diffusion v1-5 (Rombach_2022_CVPR). PyTorch graph sizes range from $1k$ to $5k$ operators.
@@ 6.2. System microbenchmarks and error analysis
Figure 8. Microbenchmarks on BERT-large dispute game: vary split size NN and dispute round index. Left: avg dispute rounds decrease roughly as $O(\log_{N}|V|)$. Mid-left: avg dispute time drops sharply from small $N$ and plateaus for $N\!\geq\!8$. Mid-right: avg Merkle proof checks for weights/graph fall as NN grows. Right: per-round substep time (proposer partition vs. challenger re-execute plus selection) measured across eight perturbed operators through the model.
A set of four graphs showing the results of microbenchmarks on the BERT-large dispute game. The first graph shows that the average number of dispute rounds decreases roughly as O(log_N—V—) as the split size N increases. The second graph shows that the average dispute time drops sharply from small N and plateaus for N¿=8. The third graph shows that the average number of Merkle proof checks for weights/graph falls as N grows. The fourth graph shows the per-round substep time (proposer partition vs. challenger re-execute plus selection) measured across eight perturbed operators through the model.
@@
Implications for verification. First, search-time checks are one–three orders of magnitude tighter than worst-case theory for transformers; evasion would require perturbations that are consistently minute across many operators, aligning with the non-accumulation results in Sec. 4. Second, leaf-only verification under theoretical bounds leaves more slack: especially for LLM stacks, motivating the committee option at the leaf when tighter adjudication is desired. We subsequently confirm these implications in the attack experiments, where empirical error percentile thresholds yield $0\%$ ASR across models while worst-case bounds allow a small attack window (Sec. 4.5).
Dispute game statistics and scaling with $N$. Fig. 8 varies the partition size $N$. Avg. Dispute Rounds: increasing $N$ reduces rounds from $\sim\!11$ at N=2N{=}2 to $\sim\!3$ at $N{\geq}12$. Avg. Dispute Time: dispute time drops sharply from N=2N{=}2 to $N{\approx}6$ and then flattens; the first round dominates latency because it handles the largest subgraph. Merkle checks: Merkle proof checks (graph and weights) shrink monotonically with $N$ (roughly $2{\times}$$4{\times}$ reduction between N=2N{=}2 and $N{\geq}8$). Per-round substep latency: the rightmost panel shows the time for the proposer’s partition and the challenger’s selection at dispute round index $i$ within an $N=4$-way partition; both decay with $i$. Boxplot information is collected through eight dispute games across eight different perturbed operators through the model. Guideline and insight: $N\!\in\![8,12]$ minimizes wall-clock and coordinator verification work without creating oversized proofs.
@@ 6.3. End-to-end cost and on-chain footprint
Table 3. Forward and dispute costs across models ($N=2$). DCR is the challenger FLOP range to reach and adjudicate the leaf; Cost Ratio is DCR/forward FLOPs. Gas sums all dispute message interactions on Ethereum Holesky testnet.
@@
Deterministic execution overhead. The deterministic settings introduced in Sec. 6.1 is also set during the optimistic execution phase (Phase 1 in Sec. 2.2) to ensure the system functions correctly. We test the overhead of these settings by running Qwen-8B model on 100 WikiText-103 (merity2017pointer) input data. The experiments show that the software deterministic settings introduce a negligible overhead of  0.3% in terms of latency compared to the non-deterministic settings.
@@
Where TAO fits. When users accept tolerance-aware correctness, TAO offers orders-of-magnitude lower end-to-end latency and memory usage than zk proving for LLM-scale inference while preserving native kernels and heterogeneity. zk remains compelling when public verifiability with model secrecy is paramount; TAO is compelling when floating-point semantics and production performance are constraints.
7. Discussion
Optimization compatibility. A practical advantage of TAO is that it is largely compatible with performance optimizations, as long as they preserve IEEE-754 semantics, including kernel-level optimizations such as tiling, unrolling, warp specialization, software pipelining, kernel fusion, instruction reordering, and parallel reduction reordering—all of which may change the rounding behavior but remain within the error bounds. In contrast, TAO does not aim to validate optimizations that change mathematical semantics or introduce input-dependent approximations (e.g., fast-math semantic changes, approximate math intrinsics via different hardware paths, undeclared quantization, early-exit strategies, and data-dependent pruning). These optimizations can still be incorporated if they are made explicit in the commitment (i.e., treated as a distinct committed graph/semantic specification) and calibrated under the corresponding configuration class.
Multi-step workloads and discrete decisions (tie-break rules). TAO naturally extends to multi-step settings (decoding, diffusion sampling, and training) by layering time over the dispute game. Concretely, we can commit to a temporal Merkle tree of step states (e.g., per-token logits, sampled tokens, MoE routing decisions, diffusion latents, or training checkpoints/optimizer state), bisect first across time to the earliest offending step, and then within that step across the operator DAG to the leaf adjudication. This yields prefix finality: early steps can finalize even while later steps remain subject to challenge. In multi-step generation, small logit deviations can flip an argmax, turning continuous numerical drift into discrete step-level divergence. To make such settings well-defined under tolerance-aware verification, TAO can enforce a deterministic, pre-committed tie-break rule when competing choices fall within the accepted tolerance. One simple rule is lexicographic tie-break among candidates whose logits are within a committed margin; another is a verifiable hash-seeded rule that deterministically selects among near-ties using a seed derived from committed public data. TAO can ensure that honest executions converge on the same discrete decision even when logits are numerically close, while disputes remain localized and adjudicable.
Onboarding new configurations. A minor limitation is calibration churn: onboarding new devices, kernels, or library stacks may shift floating-point behavior outside previously committed empirical thresholds, potentially causing benign disputes. Operationally, this can be mitigated by (i) whitelisting allowable optimization classes/configuration families for a deployment and calibrating under them, and (ii) treating newly observed benign drift as an onboarding event resolved by adjudication and followed by efficient threshold updates for the new configuration class.
Limitations TAO is designed for settings where verification via re-execution is feasible, e.g., open-model deployments or permissioned/regulated environments in which authorized verifiers can reproduce execution for dispute resolution. In practice, a regulated or contract-bound middle layer can enforce both service correctness and confidentiality.
@@ 8.1. Verifiable Machine Learning
Work on verifiable ML splits into cryptographic validity proofs and optimistic approaches. Zero-knowledge systems compile models to finite-field circuits and attach succinct proofs (e.g., SafetyNets (ghodsi2017safetynets), zkML (chen2024zkml), zkGPT (qu2025zkgpt), zkllm (sun2024zkllm)), but still incurring substantial proving time/memory and often requiring quantization or restricted operators; TEE-assisted designs reduce proof cost but reintroduce trust in hardware vendors and exists performance degradation due to limited storage and compute capacity (tramerslalom; moon2025asgard; bai2025phantom; niu20223legrace). Optimistic protocols (Agatha (zheng2021agatha), opML (conway2024opml)) and refereed delegation (Verde (arun2025verde)) localize disputes on computation graphs and arbitrate on-chain, with practicality often hinging on replication or enforcing bitwise-reproducible operators (RepOps) across heterogeneous accelerators; recent systems generalize dispute games to Turing-complete programs and explore incentive mechanisms such as Proof-of-Sampling (mirkin2025arbigraph; zhang2024proof). TAO takes a complementary tack: rather than enforcing bitwise equality, it verifies results up to principled, operator-level tolerances that acknowledge nondeterminism. Unlike approaches that determinize execution (e.g., high-precision rounding logs for verifiable training or RepOps libraries (srivastava2024optimistic; arun2025verde)), TAO aims to preserve native GPU kernels and heterogeneity while still providing portable accountability.
@@ 8.2. Floating-Point Rounding Errors in Neural Networks
A growing body of work shows that ignoring floating-point effects renders verification unsound in practice: real-number proofs can certify networks that fail under IEEE-754 execution, and deployed stacks exhibit nondeterminism from non-associative reductions, kernel fusion, and scheduler variation (jia2021exploiting; szasz2025no). Measurement tools such as FPRev reverse-engineer accumulation orders across software/hardware stacks(xie2025revealing). In response, some systems try to eliminate nondeterminism, e.g., reproducible operator libraries or training with high-precision states and logged rounding decisions to enable exact replay across GPUs, but at non-trivial performance and engineering cost (arun2025verde; srivastava2024optimistic).
9. Conclusion
We introduce TAO, a tolerance-aware optimistic protocol that makes neural network computation verifiable on heterogeneous accelerators without relying on bitwise determinism or trusted hardware. TAO verifies results up to operator-specific acceptance regions by combining portable IEEE-754 bounds with tight empirical error percentile thresholds, and it localizes disputes to a single operator via a Merkle-anchored dispute game followed by a low-cost leaf check. This design preserves native kernels and performance while providing vendor-agnostic accountability; our implementation and attack study demonstrate the robustness and practicality of tolerance-aware verification for real-world ML workloads.